The tables below show the demand curves for the three consumers in a market (D1, D2, and D3) and the supply curves for the three producers in the market (S1, S2, and S3). First, solve for the market demand curve and the market supply curve by finding the market demand and supply at the prices of $10, $15, $20, and $25. Then find the equilibrium price and quantity in the market.

Price D1 D2 D3
10 30 8 27
15 20 4 9
25 0 1 1
Price S1 S2 S3
10 1 2 1
15 2 4 3
20 3 6 6
25 4 8 9
Answer #1
Price D1 D2 D3 Market demand = D1 + D2 + D3
10 30 8 27 30 + 8 + 27 = 65
15 20 4 9 20 + 4 +9 = 33
25 0 1 1 0 + 1 + 1 = 2
Price S1 S2 S3 Market supply = S1+S2+S3
10 1 2 1 1+2+1 = 4
15 2 4 3 2+4+3 = 9
20 3 6 6 3+6+6 = 15
25 4 8 9 4+8+9 = 21

Demand information for $20 is not given. Equilibrium occurs at the point where Market demand = Market supply. So, it might occur at equilibrium price, P = $20 and thus equilibrium quantity will be 15.